# AnDB: Breaking Boundaries with an AI-Native Database for Universal Semantic Analysis

### v1

@@ 2.2. Query Optimization
By the cost models, AnDB can select an execution plan that minimizes financial cost while achieving the desired level of accuracy.
@@ 2.3. Execution Engine
LLM invocations primarily rely on API integration. However, if local deployment is required, there are numerous of significant potential for optimization in areas such as resource scheduling, GPU utilization, and in-memory data caching. Additionally, AnDB has a lightweight knowledge graph module, inspired by GraphRAG  (Edge et al., 2024), to aid in reasoning, as graphs effectively represent entity relationships. The execution engine can retrieve structured, semi-structured, and unstructured data from disk. A key optimization involves indexing unstructured files, such as extracting keywords from images using multimodal models (e.g., FUYU-8b (Li et al., 2023)) and preprocessing documents (e.g., chunking) before storage because performing these tasks during execution would be inefficient.
@@ 3. DEMONSTRATION OVERVIEW
SELECT area, SEM_GROUP("Count the number of publications in each area")
FROM TABULAR(PROMPT("technical area") as area FROM FILE("neurips_2024.txt"))
@@
Acknowledgements.
We would like to express our gratitude to David Anugraha and Sona Ghotra for providing us with early discussions and code contributions.

### v2

@@ Abstract.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 2.2. Query Optimization
By using cost models, AnDB can select an execution plan that minimizes financial costs while achieving the desired level of accuracy.
@@ 2.3. Execution Engine
LLM invocations primarily rely on API integration. However, if local deployment is required, there is a lot of significant potential for optimization in areas such as resource scheduling, GPU utilization, and in-memory data caching. Additionally, AnDB has a lightweight knowledge graph module, inspired by GraphRAG  (Edge et al., 2024), to aid in reasoning, as graphs effectively represent entity relationships. The execution engine can retrieve structured, semi-structured, and unstructured data from disk. A key optimization involves indexing unstructured files, such as extracting keywords from images using multimodal models (e.g., FUYU-8b (Li et al., 2023)) and preprocessing documents (e.g., chunking) before storage because performing these tasks during execution would be inefficient.
@@ 3. DEMONSTRATION OVERVIEW
SELECT count(area), SEM_GROUP(title, "Area of publications", 5) /* 5 is optional, which means divided into five groups */ as area
FROM TABULAR(PROMPT("title of the paper") as title FROM FILE("neurips_2024.txt"))